Absolute value problem
When a=-4, b=-7, and c=9
|a + b + c|-2 =

Found 2 solutions by JBarnum, stanbon:
Answer by JBarnum(2146)

|a + b + c|-2 =
|-4 + -7 + 9|-2 =
|-11 + 9|-2 =
|-2|-2 =
2-2=0

Answer by stanbon(75887)

Absolute value problem
When a=-4, b=-7, and c=9
|a + b + c|-2 =
----
|-4 + -7 + 9}-2
----
= |-2|-2
---
= 2-2
=0
=========
Cheers,
Stan H.